What is the maximum strength of the magnetic field (T) in an electromagnetic wave that has...

  1. What is the maximum strength of the magnetic field (T) in an electromagnetic wave that has a maximum electric field strength of 1.00 x 103 V/m?
  1. 3.33 x 10-6
  2. 1.01 x 10-6
  3. 5.98 x 10-5
  4. 9.00 x 10-4
Answer #1

a. 3.33 x 10 -6

*********************

Maximum electric and magnetic field strength are related by :

B_{max} = \frac{E_{max}}{c}

Where c is speed of light = 3 x 10 8 m/s

Thus

B_{max} = \frac{10^3}{3\times10^8} T = 3.33\times10^{-6}\;T